OpenStudy (anonymous):

find the limit (1-e^x)^(2/x) Limit as x approaches 0

OpenStudy (mathmate):

To find L=Lim (1-e^x)^(2/x) as x->0 Take log: L'=lim (2/x)*log(1-e^x) as x->0 =lim 2log(1-e^x)/x Apply de l'hôpital's rule =lim -2e^x/(1-e^x)*1 as x->0 =inf => L=e^L'=inf